# Theory of In-Plane-Magnetic-Field-Dependent Excitonic Spectra in Atomically Thin Semiconductors

### v1

@@ Abstract
The linear absorption spectrum of excitons in TMDC monolayers under the influence of an in-plane magnetic field is theoretically studied. We demonstrate that in-plane magnetic fields induce a hybridization between spin-bright and spin-dark exciton transitions, resulting in a brightening of spin-dark excitons. We analytically investigate spectral features including resonance energy shifts, broadening and amplitudes ratios. In particular, for a MoSe2 monolayer with radiatively-limited linewidth, we find a complex interplay of dark-bright splitting and linewidth difference of both involved spin-bright and spin-dark excitons.
@@ I Introduction
In this paper, we employ a thorough analytical approach to the effects of in-plane magnetic fields on the energetically lowest excitons in atomically thin semiconductors. We present full analytical expressions for the optical absorption, hybridized energies, linewidths and amplitudes including dissipative processes such as radiative decay or phonon-assisted decay. The analytical formulas can be easily used to analyze experimental spectra. We distinguish between two classes of materials: spin-dark materials such as MoS2, WS2 and WSe2 with lowest-lying excitonic states of unequal spin, which usually show an energetic separation between the spin-dark and spin-bright state exceeding 10 meV [20, 31, 32], and spin-bright materials such as MoSe2, which exhibit lowest-lying excitonic states of equal spin and a dark-bright splitting close to 1 meV [19, 20]. The paper is organized as follows: First, we develop the equations of motion for the spin-bright and spin-dark excitonic dipole (Sec. II) and their solutions (Sec. III) under the influence of an in-plane magnetic field, then, we calculate the absorption spectra of two example TMDC monolayers of the two different classes, MoSe2 and MoS2 (Sec. IV), and, at last, we analytically analyze the occurring magnetic-field-dependent spectral features (Sec. IV.1Sec. IV.3).
@@ II Equations of Motion
Here, $P^{\,\xi,s,\xi,s}_{\nu}$ is the excitonic transition [34]:
$P^{\,\xi,s_{1},s_{2}}_{\nu}=\sum_{\mathbf{q}}\varphi_{\nu,\textbf{q}}^{*,\xi,s_{1},s_{2}}\langle\hat{v}^{\dagger,\xi,s_{1}}_{\mathbf{q}}\hat{c}^{\xi,s_{2}}_{\mathbf{q}}\rangle,$ (2)
@@
Here, quadratic contributions are neglected as the effect is already small in the linear regime and, to the best of our knowledge, no diamagnetic shifts have been measured yet with an in-plane geometry for TMDC monolayers. The in-plane interaction strength is governed by the excitonic magnetic matrix elements:
@@
where $\mu_{B}=\frac{e\hbar}{2m_{0}}$ is the Bohr magneton and $B_{\parallel}$ is the magnetic field strength. The $g$-factor can approximately be described by the free-electron $g$-factor $g_{e}\approx 2$ [5, 19, 22]. Note that $\overline{s}$ corresponds to the opposite spin of $s$, i.e., if $s=\,\uparrow$, then $\overline{s}=\,\downarrow$ and vice versa, which is caused by the action of an in-plane magnetic field on the spin wave functions. Therefore, the magnetic Hamiltonian is non-diagonal with respect to the out-of-plane spin states of the electrons and holes of the corresponding excitons. The consequences of this situation for linear optical spectra are the main focus of this paper. The magnetic matrix elements in Eq. 6 determine the spin-coupling selection rules: Due to the spatial homogeneity of the in-plane magnetic field $B_{\parallel}$, a coupling of angle-independent $s$-orbital states with angle-dependent $p$-orbital states does not occur. Moreover, a coupling between the $1s$-exciton and higher-lying $ns$-excitons quickly fades as $n$ increases due to a decreasing overlap of the corresponding exciton wave functions. Note that the former is a characteristic property of the interaction of a spatially homogeneous in-plane magnetic field with the spin angular momentum, as excitonic spin flips via spin-orbit interaction within spatially homogeneous [29, 36] or spatially inhomogeneous [37] out-of-plane electric fields, i.e., Rashba interaction [28], behave differently .
@@
which consists of a non-radiative dephasing $\gamma_{\text{nrad}}^{\xi,s_{1},s_{2}}$ with a contribution due to exciton-phonon interaction, which can be calculated [38], or due to disorder [39] or strain [40], which are introduced phenomenologically, and a radiative contribution $\gamma_{\text{rad},\nu}^{\xi,s_{1},s_{2}}=\frac{\varepsilon_{x,\nu}^{\xi,s_{1},s_{2}}}{\hbar^{2}2\epsilon_{0}c_{0}n_{\text{refr}}}|d_{\xi,s_{1}}^{\,c,v}|^{2}\delta_{s_{1},s_{2}}$ due to reradiation [33] with vacuum permittivity $\epsilon_{0}$, speed of light $c_{0}$ and mean refractive index $n_{\text{refr}}=\frac{1}{2}(\sqrt{\epsilon_{1}}+\sqrt{\epsilon_{2}})$ of the homogeneous media surrounding the TMDC monolayer from both sides. Note, that we assume $\epsilon_{1}=\epsilon_{2}$ throughout this work. The second line describes the coupling to an optical field via the excitonic Rabi frequency:
@@ III Hybridized Excitonic States
Without loss of generality, in the following, we restrict ourselves to right-handed polarization $\sigma=\sigma_{+}$ that can optically excite electron-hole pairs at the $K$-valley only. Since we work in the regime of linear optics, our description is equally valid for left-handed or linear optical excitation and we can neglect intervalley coupling mechanisms, which are mostly relevant in nonlinear dynamics [41, 42, 43, 44, 45, 46, 47] or states with non-vanishing center-of-mass momenta Q [48, 49]. Moreover, we restrict the description to the lowest-lying $1s$-excitons of the A-series, where we neglect A/B-coupling, since it is strongly suppressed due to the large valence band splitting of several hundreds of meV [4]. The remaining two equations of motion from Eq. 7 can be solved analytically in frequency space by first eliminating the spin-dark ($d$) excitonic transition in the emerging analytical expression of the spin-bright ($b$) excitonic transition. The spin-bright transition $P^{b}(\omega)=P_{1s}^{\xi,s,s}$, which couples directly to the Maxwell field via Eq. 1 and contains – due to spin mixing – also spectral signatures of the dark transition, then reads:
@@
We note that due to the explicit inclusion of different linewidths of the involved spin-bright and spin-dark excitons, which directly impact the spin hybridization, our excitonic description resembles a non-Hermitian treatment to the optical response of atomically thin semiconductors [50]. These non-Hermiticy-resembling effects are most pronounced, if energy splitting, linewidth and linewidth differences between the involved resonances are all of similar magnitude. Such a regime can occur in, e.g., disorder-/defect-free h-BN-encapsulated MoSe2 monolayers at cryogenic temperatures with a dark-bright energy splitting $\Delta$ of 1.5 meV and an almost radiatively-limited linewidth $\hbar\gamma$ of the spin-bright exciton of approximately 111.51.5\,meV.
@@
$\lim_{B_{\parallel}\mapsto\infty}P^{\mathcal{S}}_{B_{\parallel}}=\lim_{B_{\parallel}\mapsto\infty}P^{\overline{\mathcal{S}}}_{B_{\parallel}}=0.5,$ (20)
for very large magnetic field strengths. However, this regime clearly exceeds currently experimentally available field strengths on the order of 100 T [12]. In Eq. 18 and Fig. 2, additional imaginary components appear because we do not neglect dissipation throughout our whole analysis.
Figure 2: Real- and imaginary part of mixing coefficients $P_{B_{\parallel}}^{\mathcal{S}}$ for $\nu=1s$ under influence of the in-plane magnetic field for MoSe2 ($\hbar\gamma_{\text{nrad}}=0.5$ meV).
Note that because of the varying energy ordering of spin-bright and spin-dark excitons in the respective sample of interest, the assignment of the $\mathcal{S}$-states to specific spin-bright or spin-dark states can change.
@@ IV Linear Absorption Spectrum
that are computed with the macroscopic polarization $P^{\sigma}(t)$ from Eq. 1. By performing a partial fraction decomposition again and, with the help of Eq. 11, the absorption can be expressed explicitly as:
@@
$\begin{split}\alpha(\omega)=\sum_{\mathcal{S}}&\frac{\mathcal{F}^{\mathcal{S}}-\mathcal{G}^{\mathcal{S}}\hbar\omega}{(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega)^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}},\end{split}$ (24)
which is a superposition of the two peak signals modulated by the amplitudes $\mathcal{F}_{\mathcal{S}},\mathcal{G}_{\mathcal{S}}$, cf. App. A. The amplitudes $\mathcal{F}_{\mathcal{S}}$ correspond to symmetric Lorentzian signals, which are modified by the amplitudes $\mathcal{G}_{\mathcal{S}}$ scaling linearly with the energy $\hbar\omega$. The latter are interference effects, which occur as the two hybridized excitonic frequencies come energetically close, causing an asymmetric deformation of the spectrum.
@@
We depict in Fig. 3 the linear absorption spectra Eq. 24 of monolayers MoSe2 (left) and MoS2 (right) encapsulated in h-BN for three in-plane magnetic field-strengths around the spin-bright 1s-orbital state. From top to bottom, we vary the non-radiative linewidth, which describes the regime of a radiatively-limited linewidth up to a regime, where phonons gain a dominating influence (100 K) [38].
@@
Figure 3: Linear absorption spectrum under influence of an in-plane magnetic field for $\nu=1s$ with different field-strengths for the material MoSe2 in (a)–(c) and MoS2 in (d)–(f) encapsulated in h-BN with $\sigma_{+}$-polarised light for increasing non-radiative linewidths $\gamma_{\text{nrad}}$.
@@
To obtain Fig. 3, we adjusted the excitonic energies according to recent PL experiments with dark-bright splitting of 1.45 meV in the optically bright MoSe2 monolayer [19, 20] and dark-bright splitting of 14 meV in the optically dark MoS2 monolayer [20]. We calculated the Coulomb enhancement $\sum_{\textbf{q}}\varphi_{1s,\textbf{q}}^{*\,K,\uparrow,\uparrow}$ in Eq. 9 mediated by the excitonic wave functions as solutions of the corresponding Wannier equations and the transitions dipole moments with $\mathbf{k}\cdot\mathbf{p}$-parameters according to Refs. [35, 4]. All parameters are provided in Tab. 1 in the appendix.
@@
In the absorption of a MoSe2 monolayer in Fig. 3 at increasing in-plane magnetic field strengths, the initially spin-dark resonance brightens, i.e., gains an oscillator strength. Simultaneously, both excitonic resonances shift in the opposite direction, i.e., an anti-crossing occurs. This has also been observed in PL [20, 52, 19, 53, 54, 21]. In the MoS2 monolayer, the same behavior occurs, but strongly suppressed due to the much larger dark-bright splitting of 14.5 meV compared to 1.45 meV in the MoSe2 monolayer. The brightening of initially spin-dark excitons is most pronounced in the radiatively-limited case of the linewidth, where a strong signal at the spin-dark resonance already appears at 15 T. At a slightly increased non-radiative linewidth of $\gamma_{\text{nrad}}=0.5$ meV, a clear spin-dark resonance appears at 30 T and a shoulder can be observed at 15 T in a MoSe2 monolayer, while, in a MoS2 monolayer, the initial spin-dark resonance is barely detectable even at 30 T. At comparably large non-radiative linewidths of $\gamma_{\text{nrad}}=5\,$meV, both resonances can no longer be separated and appear as one single, broadened peak.
Also, difficult to detect by eye, but present, is a redistribution of the linewidths in an increasing magnetic field, which is modelled by Eq. 13.
@@
These three magnetic-field-dependent features – the peak position, the linewidth and the behaviour of the amplitude – are discussed in detail in the following.

### v2

@@ Abstract
The linear absorption spectrum of excitons in TMDC monolayers under the influence of an in-plane magnetic field is theoretically studied. We demonstrate that in-plane magnetic fields induce a hybridization between spin-bright and spin-dark exciton transitions, resulting in a brightening of spin-dark excitons. We analytically investigate spectral features including resonance energy shifts, broadening and amplitudes ratios. In particular, for a MoSe2 monolayer with linewidths dominated by reradation, we find a complex interplay of dark-bright splitting and linewidth difference of both involved spin-bright and spin-dark excitons.
@@ I Introduction
In this paper, we employ a thorough analytical approach to the effects of in-plane magnetic fields on the energetically lowest excitons in atomically thin semiconductors. We present full analytical expressions for the optical absorption, hybridized energies, linewidths and amplitudes including dissipative processes such as radiative decay or phonon-assisted decay. The analytical formulas can be easily used to analyze experimental spectra. We distinguish between two classes of materials: spin-dark materials – such as MoS2, WS2 and WSe2 – , where the spin-dark state lies energetically below the spin-bright state; in this case, their energetic separation exceeds usually 10 meV [50, 10, 45]. On the other hand, spin-bright materials – such as MoSe2 – exhibit a spin-bright state that lies energetically below the spin-dark state; these states are energetically separated by roughly 1 meV [39, 50]. The paper is organized as follows: First, we develop the equations of motion for the spin-bright and spin-dark excitonic dipole (Sec. II) and their solutions (Sec. III) under the influence of an in-plane magnetic field, then, we calculate the absorption spectra of two example TMDC monolayers of the two different classes, MoSe2 and MoS2 (Sec. IV), and, at last, we analytically analyze the occurring magnetic-field-dependent spectral features (Sec. IV.1Sec. IV.3).
@@ II Equations of Motion
Here, $P^{\,\xi,s,s}_{\nu}$ is the excitonic transition [26]:
$P^{\,\xi,s_{1},s_{2}}_{\nu}=\sum_{\mathbf{q}}\varphi_{\nu,\textbf{q}}^{*\,\xi,s_{1},s_{2}}\langle\hat{v}^{\dagger,\xi,s_{1}}_{\mathbf{q}}\hat{c}^{\xi,s_{2}}_{\mathbf{q}}\rangle,$ (2)
@@
We neglect diamagnetic contributions [60, 76] for the molybdenum-based materials for an in-plane geometry because they are sufficiently small (roughly 0.03 meV at 30 T according to our computations) such that they are not of major importance – even though it causes subtle asymmetries for the level repulsion in Fig. 4 – and it is beyond the scope of this paper. The in-plane interaction strength is governed by the excitonic magnetic matrix elements:
@@
where $\mu_{B}=\frac{e\hbar}{2m_{0}}$ is the Bohr magneton and $B_{\parallel}$ is the magnetic field strength. The $g$-factor can approximately be described by the free-electron $g$-factor $g_{e}\approx 2$ [33, 39, 25] and are not valley-dependent or influenced by Berry-curvature contributions due to the in-plane geometry [59, 71]. Note that $\overline{s}$ corresponds to the opposite spin of $s$, i.e., if $s=\,\uparrow$, then $\overline{s}=\,\downarrow$ and vice versa, which is caused by the action of an in-plane magnetic field on the spin wave functions. Therefore, the magnetic Hamiltonian is non-diagonal with respect to the out-of-plane spin states of the electrons and holes of the corresponding excitons. The consequences of this situation for linear optical spectra are the main focus of this paper. The magnetic matrix elements in Eq. 6 determine the spin-coupling selection rules: Due to the spatial homogeneity of the in-plane magnetic field $B_{\parallel}$, a coupling of angle-independent $s$-orbital states with angle-dependent $p$-orbital states does not occur. Moreover, a coupling between the $1s$-exciton and higher-lying $ns$-excitons quickly fades as $n$ increases due to a decreasing overlap of the corresponding exciton wave functions. Note that the former is a characteristic property of the interaction of a spatially homogeneous in-plane magnetic field with the spin angular momentum, as excitonic spin flips via spin-orbit interaction within spatially homogeneous [8, 43] or spatially inhomogeneous [42] out-of-plane electric fields, i.e., Rashba interaction [70], behave differently. Lastly, we note that due to the spin-orbit coupling, which is effectively incorporated in the excitonic Hamiltonian $\hat{H}_{\text{X-0}}$ and diagonal in the out-of-plane spin basis [31, 32, 30], the excitonic Hamiltonian does not contain any contributions from in-plane spins, which leads to an in-plane rotational spin invariance. Additionally, the spin part is not affected by the lattice geometry because the Hamiltonian factorizes the Bloch part – which encodes the geometry – and the spin part. As of such, different in-plane field-orientations (e.g. zig-zag or armchair) yield equivalent results.
@@
which consists of a non-radiative dephasing $\gamma_{\text{nrad}}^{\xi,s_{1},s_{2}}$ with a contribution due to exciton-phonon interaction, which can be calculated and differs for respective monolayer materials due to different excitonic landscapes [55], or due to disorder [62] and strain [28], which are introduced phenomenologically, and a radiative contribution $\gamma_{\text{rad},\nu}^{\xi,s_{1},s_{2}}=\frac{\varepsilon_{x,\nu}^{\xi,s_{1},s_{2}}}{\hbar^{2}2\epsilon_{0}c_{0}n_{\text{refr}}}|d_{\xi,s_{1}}^{\,c,v}|^{2}\delta_{s_{1},s_{2}}$ due to reradiation [29] with vacuum permittivity $\epsilon_{0}$, speed of light $c_{0}$ and mean refractive index $n_{\text{refr}}=\frac{1}{2}(\sqrt{\epsilon_{1}}+\sqrt{\epsilon_{2}})$ of the homogeneous media surrounding the TMDC monolayer from both sides. Note that we assume $\epsilon_{1}=\epsilon_{2}$ throughout this work. The second line describes the coupling to an optical field via the excitonic Rabi frequency:
@@ III Hybridized Excitonic States
Without loss of generality, in the following, we restrict ourselves to right-handed polarization $\sigma=\sigma_{+}$ that can optically excite electron-hole pairs at the $K$-valley only. Since we work in the regime of linear optics, our description is equally valid for left-handed or linear optical excitation and we can neglect intervalley coupling mechanisms, which are mostly relevant in nonlinear dynamics [57, 27, 11, 3, 54, 35, 15] or states with non-vanishing center-of-mass momenta Q [49, 72]. Moreover, we restrict the description to the lowest-lying $1s$-excitons of the A-series, where we neglect A/B-coupling, since it is strongly suppressed due to the large valence band splitting of several hundreds of meV [30]. The remaining two equations of motion from Eq. 7 can be solved analytically in frequency space by first eliminating the spin-dark ($d$) excitonic transition in the emerging analytical expression of the spin-bright ($b$) excitonic transition. The spin-bright transition $P^{b}(\omega)=P_{1s}^{K,\uparrow,\uparrow}(\omega)$, which couples directly to the Maxwell field via Eq. 1 and contains – due to spin mixing – also spectral signatures of the dark transition, then reads:
@@
To obtain a form that reveals the dependencies of the mixing coefficient transparently – while losing the compactness of the expression – , we can insert the definitions Eqs. 12 and 13 and transform to polar coordinates; from this, we learn that the mixing coefficient does not parametrically depend on individual linewidths nor energies but only on the dark-bright splitting $\Delta$ and linewidth difference $\kappa$.
We note that due to the explicit inclusion of different linewidths of the involved spin-bright and spin-dark excitons, which directly impact the spin hybridization, our excitonic description resembles a non-Hermitian treatment to the optical response of atomically thin semiconductors [69]. These non-Hermiticy-resembling effects are most pronounced, if energy splitting, linewidth and linewidth differences between the involved resonances are all of similar magnitude. Such a regime can occur in, e.g., disorder-/defect-free h-BN-encapsulated MoSe2 monolayers at cryogenic temperatures with a dark-bright energy splitting $\Delta$ of 1.5 meV and a linewidth $\hbar\gamma$ of the spin-bright exciton of approximately 111.51.5\,meV.
@@
$\lim_{B_{\parallel}\mapsto\infty}P^{\mathcal{S}}_{B_{\parallel}}=\lim_{B_{\parallel}\mapsto\infty}P^{\overline{\mathcal{S}}}_{B_{\parallel}}=0.5$ (20)
for very large magnetic field strengths. However, this regime clearly exceeds currently experimentally available field strengths on the order of 100 T [22]. Even without an in-plane magnetic field, intrinsic spin-hybridization at the $K$ valleys [12, 24] is present. However, in molybdenum-based materials, this spin-hybridization is much less than 10 %. Hence, within our model, we can safely assume an ideal spin model, since the spin-hybridization induced by in-plane magnetic fields at 30 T – cf. Fig. 2 – assumes values around 30 % that greatly exceeds the intrinsic spin-hybridization. In Eq. 18 and Fig. 2, additional imaginary components appear because we do not neglect dissipation throughout our whole analysis.
Figure 2: Real- and imaginary part of mixing coefficients $P_{B_{\parallel}}^{\mathcal{S}}$ for $\nu=1s$ under influence of the in-plane magnetic field for MoSe2 ($\kappa=0.5$ meV).
Lastly, we note that, due to varying energy ordering of spin-bright and spin-dark excitons in the respective sample of interest, the assignment of the $\mathcal{S}$-states to specific spin-bright or spin-dark states can change.
@@ IV Linear Absorption Spectrum
that are computed with the macroscopic polarization $P^{\sigma}(t)$ from Eq. 1. By performing a partial fraction decomposition again, making a close-to-resonance approximation (discussed in App. A) and with the help of Eq. 11, the absorption spectrum can be expressed explicitly as:
$\alpha(\omega)=\sum_{\mathcal{S}}\frac{\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}}{(\hbar\omega_{x,B_{\parallel}}^{\mathcal{S}}-\hbar\omega)^{2}+(\hbar\gamma_{B_{\parallel}}^{\mathcal{S}})^{2}},$ (24)
with the spin-hybridized amplitudes:
@@
$\begin{split}\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}&=-2(\hbar\gamma_{\text{rad}}^{b})^{2}|P^{\mathcal{S}}_{B_{\parallel}}|^{2}\\ &\quad+2\hbar\gamma_{B_{\parallel}}^{\mathcal{S}}\big(\hbar\gamma_{\text{rad}}^{b}\text{Re}[P^{\mathcal{S}}_{B_{\parallel}}]+2\mathcal{S}\text{Im}[\Lambda_{B_{\parallel}}]\big),\end{split}$ (25)
which is a superposition of the two peak signals with amplitudes $\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}$ that effectively incorporate interference effects: The first term on the right hand side of Eq. 25 comprises the amplitude – present without applying an in-plane magnetic field – that is redistributed among the two spin-hybridized resonances; the second term originates from interference between the transmitted field and the incident optical field; the third term that scales with $\Lambda_{B_{\parallel}}$ (defined in App. A) arises from interference between the two spin-hybridized resonances. This formula provides a motivation for experimentalists to fit their spectra with a pair or Lorentzians. Its derivation is discussed in App. A.
@@
We depict in Fig. 3 the linear absorption spectra Eq. 24 of monolayers MoSe2 (left) and MoS2 (right) encapsulated in h-BN for three in-plane magnetic field-strengths around the spin-bright 1s-orbital state. From top to bottom, we vary the non-radiative linewidth $\hbar\gamma_{\text{nrad}}^{b/d}=\hbar\gamma_{\text{nrad}}$, which describes the regime of a linewidth that is dominated by reradiation up to a regime, where phonons gain a dominating influence (100 K) [55]. Additionally, we display the two individual resonance components ($\mathcal{S}=1$ and $\mathcal{S}=-1$) from Eq. 24 at 30 T. We chose an h-BN-embedding because it enables high-purity samples with comparably small linewdiths [58, 5, 1].
@@
Figure 3: Linear absorption spectrum under influence of an in-plane magnetic field for $\nu=1s$ with different field-strengths for the material MoSe2 in (a)–(c) and MoS2 in (d)–(f) encapsulated in h-BN with $\sigma_{+}$-polarised light for increasing non-radiative linewidths $\hbar\gamma_{\text{nrad}}$ as tuning parameter
@@
To obtain Fig. 3, we adjusted the excitonic energies according to recent PL experiments with dark-bright splitting of 1.45 meV in the optically bright MoSe2 monolayer [39, 50] and dark-bright splitting of 14.5 meV in the optically dark MoS2 monolayer [50]. We calculated the Coulomb enhancement $\sum_{\textbf{q}}\varphi_{1s,\textbf{q}}^{*\,K,\uparrow,\uparrow}$ in Eq. 9 mediated by the excitonic wave functions as solutions of the corresponding Wannier equations and the transitions dipole moments with $\mathbf{k}\cdot\mathbf{p}$-parameters according to Refs. [44, 30]. All parameters are provided in Tab. 2 in the appendix.
@@
In the absorption of a MoSe2 monolayer in Fig. 3 at increasing in-plane magnetic field strengths, the initially spin-dark resonance brightens, i.e., gains an oscillator strength. Simultaneously, both excitonic resonances shift in the opposite direction, i.e., an anti-crossing occurs. This has also been observed in PL [50, 19, 39, 75, 46, 73]. In the MoS2 monolayer, the same behavior occurs, but strongly suppressed due to the much larger dark-bright splitting of 14.5 meV compared to 1.45 meV in the MoSe2 monolayer. The brightening of initially spin-dark excitons is most pronounced with a negligible non-radiative linewidth compared to the radiative linewidth – cf. Fig. 3(a) –, where a strong signal at the spin-dark resonance already appears at 15 T. At a slightly increased non-radiative linewidth of $\hbar\gamma_{\text{nrad}}=0.5$ meV, a clear spin-dark resonance appears at 30 T and a shoulder can be observed at 15 T in a MoSe2 monolayer, while, in a MoS2 monolayer, the initial spin-dark resonance is barely detectable even at 30 T. At comparably large non-radiative linewidths of $\hbar\gamma_{\text{nrad}}=5\,$meV, both resonances can no longer be separated and appear as one single, broadened peak. Therefore, we require small dark-bright splitting for strong spin-hybridization and sufficiently non-overlapping resonances such that the spin-hybridization remains optically detectable.
We observe an enhanced amplitude – even without in-plane magnetic field – when raising the non-radiative linwidth from $\hbar\gamma_{\text{nrad}}=0.1$ meV to $\hbar\gamma_{\text{nrad}}=0.5$. This is due to stronger interference between the initial optical field and the spin-hybridized resonances (second term from Eq. 25), which leads to enhanced absorption. The amplitude drops for $\hbar\gamma_{\text{nrad}}=5.0$ meV because a further increase of amplitude is canceled by broadening. We note that normalized values (w.r.t. spin-bright amplitude without magnetic field and corresponding non-radiative linewidth) of the spin-hybridized amplitude $\mathcal{A}^{\mathcal{S}}_{B_{\parallel}}$ show a negligible dependence on the non-radiative linewidth. This implies that, while the non-radiative linewidth governs the absolute magnitudes of the spin-hybridized amplitudes $\mathcal{A}_{B_{\parallel}}^{\mathcal{S}}$, it does not affect the global spectral form. Also, difficult to detect by eye, but present, is a redistribution of the linewidths in an increasing magnetic field, which is modelled by Eq. 13. Also, difficult to detect by eye, but present, is a redistribution of the linewidths in an increasing magnetic field, which is modelled by Eq. 13.
@@
These three magnetic-field-dependent features – the peak position, the linewidth and the behavior of the amplitude – are discussed in detail in the following.